# Quantifying Fairness in LLMs Beyond Tokens: A Semantic and Statistical Perspective

### v2

@@ Abstract
Large Language Models (LLMs) often generate responses with inherent biases, undermining their reliability in real-world applications. Existing evaluation methods often overlook biases in long-form responses and the intrinsic variability of LLM outputs. To address these challenges, we propose FiSCo (Fine-grained Semantic Computation), a novel statistical framework to evaluate group-level fairness in LLMs by detecting subtle semantic differences in long-form responses across demographic groups. Unlike prior work focusing on sentiment or token-level comparisons, FiSCo goes beyond surface-level analysis by operating at the claim level, leveraging entailment checks to assess the consistency of meaning across responses. We decompose model outputs into semantically distinct claims and apply statistical hypothesis testing to compare inter- and intra-group similarities, enabling robust detection of subtle biases. We formalize a new group counterfactual fairness definition and validate FiSCo on both synthetic and human-annotated datasets spanning gender, race, and age. Experiments show that FiSCo more reliably identifies nuanced biases while reducing the impact of stochastic LLM variability, outperforming various evaluation metrics.
@@ 1 Introduction
Several methods have been developed to detect and measure potential biases in LLMs. Demographic representation evaluation methods (Brown et al., 2020; Liang et al., 2022; Mattern et al., 2022; Smith et al., 2022b; Abid et al., 2021; Barikeri et al., 2021) confine the definition of bias to a narrow and commonly used vocabulary. Consequently, these methods fail to capture group-level differences across demographic attributes. Similarly, counterfactual fairness methods (Smith et al., 2022b; Barikeri et al., 2021; Nadeem et al., 2020) typically focus on certain elements in the model’s response rather than the overall response. As a result, they struggle to detect semantic differences in long-form outputs. While some recent efforts (Dwivedi-Yu et al., 2024) recognize the difficulty of capturing subtle biases and attempt to measure LLM fairness from a broader perspective, existing bias benchmarks are limited in length (fewer than 400 characters; see Appendix I for analysis), while GPT-4 produces responses averaging more than 600 characters (Chen et al., 2023).
To address these challenges, we propose FiSCo (Fine-grained Semantic Computation), a novel statistical framework for evaluating group-level fairness in LLMs by detecting subtle semantic differences in long-form responses across demographic groups (see Figure 1). In contrast to prior methods that rely on predefined bias categories or simplistic text substitutions, our approach captures deeper, group-level disparities in model behavior. Specifically, we extend traditional group fairness concepts to settings where responses are long and variable in structure. FiSCo goes beyond token-level or surface-level comparisons by decomposing each response into semantically distinct claims, and applies state-of-the-art reference checks at the claim level (Hu et al., 2024) to evaluate alignment across responses. We then statistically assess both inter-group and intra-group variability to uncover meaningful differences between demographic groups (see Figure 2). This design enables the detection of subtle biases that would otherwise be obscured by response variability or surface-level text similarity. Furthermore, we leverage statistical hypothesis testing to robustly assess whether observed differences are significant, avoiding the pitfalls of manual prompt manipulation or handcrafted metrics that may overlook real-world patterns. By focusing on semantic content rather than superficial text features, our method enables interpretable and scalable fairness evaluation. To validate the effectiveness of our framework, we construct a comprehensive dataset spanning multiple fairness axes, including gender, race, and age. Experiments demonstrate that FiSCo consistently outperforms existing metrics in detecting nuanced biases with greater stability and alignment to human judgment. The main contributions of this paper are summarized as follows:
@@
  • A new large-scale benchmark dataset designed for long-form fairness evaluation, spanning major demographic dimensions (gender, race, and age), enriched with high-quality human annotations and realistic prompt templates, supporting comprehensive, reproducible, and multi-axis fairness assessments across multiple state-of-the-art LLMs.
@@ 2.1 Fairness Evaluation in LLMs
Figure 2: The pipeline for detecting bias in LLMs begins with prompt group generation, where base questions are modified to represent different demographic groups (e.g., male vs. female). Next, response generation produces LLM outputs for each group version of the prompt. After that, we perform claim extraction & entailment labeling, decomposing responses into semantically distinct claims and applying entailment checks between response pairs. Then, a statistical comparison is conducted by computing similarity scores between intra-group (e.g., male–male and female–female) and inter-group pairs (e.g., male–female), using Welch’s t-test to check for significant differences. If inter-group differences are statistically greater than intra-group differences, it suggests that the model exhibits bias for that question.
@@ 2.2 Comparison with Existing Datasets
BBQ (Parrish et al., 2022) Question answering $<$20 tokens Race, Gender, Age, etc. Ambiguous QA format
HolisticBias (Smith et al., 2022a) Prompt continuation $<$20 tokens 13 social groups Template-based completion
FiSCo (Ours) Long-form generation ~600 chars Gender, Race, Age Open-ended, human-validated
@@
Symbol Description
$\mathcal{M}$ A large language model (LLM)
$\theta$ Parameters of the LLM
$P_{X}$ Population distribution of prompts related to topic $X$
$X$ A topic or base prompt designed to elicit potential bias
$\epsilon$ Tolerance level for fairness conditions
$B(\cdot)$ A statistical metric applied to the model’s output
$T(\cdot,\cdot)$ Invariance metric for group-level differences
$k$ Number of questions per group
$C_{1}=\{c_{1}^{(1)},\ldots,c_{1}^{(m)}\}$ Claims extracted from response $r_{1}$
$C_{2}=\{c_{2}^{(1)},\ldots,c_{2}^{(n)}\}$ Claims extracted from response $r_{2}$
$\alpha$ Weight for claims labeled as Entailment
$\beta$ Weight for claims labeled as Neutral
$\gamma$ Weight for claims labeled as Contradiction
$C_{E}$ Count of claims labeled as Entailment
$C_{N}$ Count of claims labeled as Neutral
$C_{C}$ Count of claims labeled as Contradiction
Table 2: List of symbols used in the paper.
@@ 3.4 Fine-grained Response Similarity
3.4.1 Claim Extraction and Entailment Checking
Our approach begins with extracting individual claims from each response. For any two responses $r_{1}$ and $r_{2}$, we first identify the set of claims $C_{1}=\{c_{1}^{(1)},\ldots,c_{1}^{(m)}\}$ from $r_{1}$ and $C_{2}=\{c_{2}^{(1)},\ldots,c_{2}^{(n)}\}$ from $r_{2}$. We then perform a bidirectional semantic entailment check. For each claim $c_{1}^{(i)}$ in $C_{1}$, we determine whether it can be entailed by $r_{2}$, and vice versa for each claim in $C_{2}$ with respect to $r_{1}$. This process results in labeling each claim as one of three categories.
@@ 3.4.1 Claim Extraction and Entailment Checking
Having multiple claims per response helps our method capture subtle differences between responses. By comparing claims, FiSCo is able to detect fine-grained semantic differences between responses.
3.4.2 Similarity Scoring
To quantify the similarity between responses, we assign scores in the range of $[0,1]$ to each label type: $\alpha$ for Entailment, $\beta$ for Neutral, and $\gamma$ for Contradiction, where typically $\alpha\geq\beta\geq\gamma$, to reflect that entailment indicates greater similarity than a neutral relationship, which in turn indicates higher similarity than contradiction.
@@ 3.4.2 Similarity Scoring
Initially, we set both $\beta$ and $\gamma$ to zero, simplifying our similarity measure to focus exclusively on entailment. This reflects our goal of identifying shared information between responses without penalizing for contradictions or neutrality in the initial analysis phase. For an ablation study varying the value of $\beta$, see Appendix E.
@@ 3.5 The Proposed FiSCo Score
With our similarity score function $\mathcal{S}(r_{1},r_{2})$, we can quantitatively assess the fairness of the LLM’s responses across different demographic groups. We do this by comparing inter-group and intra-group similarities. Specifically, for each pair of groups $(G^{\prime},G^{\prime\prime})$, we compute (i) Inter-group similarities: $\mathcal{S}_{\text{inter}}=[\mathcal{S}(r_{i}^{{}^{\prime}},r_{j}^{{}^{\prime\prime}})|i,j\in\{1,\ldots,k\}]$, resulting in $k^{2}$ values, and (ii) Intra-group similarities: $\mathcal{S}_{\text{intra}}=[\mathcal{S}(r_{i}^{{}^{\prime}},r_{j}^{{}^{\prime}})|i,j\in{1,\ldots,k},i<j]+[\mathcal{S}(r_{i}^{{}^{\prime\prime}},r_{j}^{{}^{\prime\prime}})|i,j\in{1,\ldots,k},i<j]$, resulting in $k(k-1)$ values.
@@
Using this t-statistic and the degrees of freedom, we calculate the p-value to determine whether the observed differences are statistically significant, with a significance level of $0.05$. Thus, it outputs a binary score. Statistical significance allows FiSCo to avoid inaccurate assessments caused by stochastic variability in LLM responses. FiSCo offers a quantifiable signal of differential treatment in generated outputs across social groups, under controlled conditions.
@@ 4.1 Data Creation
To reflect real-world applicability, each generated template was reviewed by five human annotators via Amazon Mechanical Turk. Annotators were asked to assess whether (1) the question is likely to be asked by a human, and (2) it is suitable for ChatGPT to answer. We retained only those templates for which both questions received a “Yes” with a confidence score above 0.8. Details of our annotation interface, labeling guidelines, and example prompts are included in Appendix B.1.
@@ 4.2 Response Collection
We collect responses from several LLMs, [^1] We use all the LLMs by invoking the Amazon Bedrock APIs (https://aws.amazon.com/bedrock/). More details are given in Appendix J. including Llama 3 70B Instruct (Meta, 2024) and Mixtral 8x7B Instruct (Jiang et al., 2024). The responses were filtered to ensure a minimum length of 30 words. To maintain data integrity, we standardized the prompt structure and balanced demographic categories based on the number of prompts per group. Incomplete or off-topic responses were excluded to ensure that the final dataset contains only high-quality, analyzable outputs. Appendix A includes the prompts and examples used for both question and response generation.
@@ 4.3 Response Evaluation
Details of our SOP, annotation interface, labeling workflows, annotator feedback, and illustrative examples are provided in Appendix B.2. Human-validated templates and fine-grained response comparisons contribute to the realism and reliability of the proposed dataset.
@@ 5 Experiments
For similarity computation, we employ the open-source tool RefChecker (Hu et al., 2024), using Llama 3.1 70B Instruct (Meta, 2024) as both the claim extractor and the checker model (see Appendix G for detailed rationale and Appendix H for ablation study). All experiments were performed on a CPU-based infrastructure, and LLM APIs were accessed through AWS Bedrock. Additional ablation studies are provided in Appendix E. In both experiments, we compare the performance of different similarity metrics or group level methods. Here, we use a paired t-test (measuring scores on the same sample across different methods) to assess whether the mean difference in agreement rates between our method and the second-best method (SentenceT5/CBleu) is statistically significant.
5.1 Similarity Metric Evaluation
Experiment Setup. To evaluate the reliability of our similarity measurement Eq. (1), we conducted experiments on two datasets: synthetic and human-annotated. For the synthetic dataset (600 pairs), we generated responses by combining pre-selected claims with pre-defined conditions that specify how the claims should relate. This design enables precise control over ground-truth by explicitly defining claim-level relationships between responses. Details of this process are provided in Appendix A. For the human-annotated dataset (383 pairs), we used the labeled dataset described in Section 4.3. Baseline methods include widely used metrics such as BERTScore and Sentence-BERT, as well as top-performing sentence embedding methods from recent benchmarks (Muennighoff et al., 2022), such as SentenceT5 embeddings.
Baseline Comparison Methods. We compare methods for measuring text similarity, spanning from traditional approaches to state-of-the-art deep learning-based techniques. Bag-of-Words (BoW) (Salton, 1983): Represents text as a collection of word frequencies, without accounting for word order. Term Frequency-Inverse Document Frequency (TF-IDF) (Salton & Buckley, 1988): Assigns higher weights to words that are more distinctive across documents. Word Mover’s Distance (WMD) (Kusner et al., 2015): Measures semantic distance by computing the minimal cumulative distance between word embeddings of two texts. SimCSE (Gao et al., 2021): A contrastive learning-based method that generates sentence embeddings for similarity comparison. BERTScore (Zhang et al., 2019): Uses contextual embeddings from BERT to compute semantic overlap between texts. Sentence-BERT (Reimers, 2019): Extends BERT to produce sentence-level embeddings, enabling efficient similarity computation via cosine distance. SentenceT5 (Ni et al., 2021): A sentence embedding model built on T5, leveraging its text-to-text paradigm to produce high-quality, scalable encodings for various natural language understanding tasks. We adopt the largest available version of SentenceT5.
@@ 5.1 Similarity Metric Evaluation
Qualitative Evaluation. We present cases where traditional baselines fail to detect biases, but our proposed method successfully identifies them (see Table 14). Text 1 and Text 2 are extracted from model responses generated to assess gender bias. In the first case, both responses contain synonymous phrases, varied sentence structures, and different contextual details. Despite these surface-level differences, the two texts are semantically aligned. Traditional methods often struggle in such scenarios due to their reliance on structural and lexical similarity. However, both BERTScore and our method capture the deeper semantic equivalence between the sentences. In a contrasting case, there is a key semantic distinction between Text 1 and Text 2, even though their wording overlaps. Text 1 recommends pursuing a master’s degree in computer science, while Text 2 advises a degree in engineering management. Although BERTScore assigns a high similarity score (0.95), this difference is critical for identifying bias. Specifically, the responses suggest that for a female software engineer, the model recommends further technical development, whereas for an equally qualified male engineer, it promotes a transition into leadership. This reflects a potential gender bias, implying that women are viewed as needing technical improvement, while men are encouraged to lead. By leveraging natural language inference (NLI) at the claim level, our proposed similarity metric can detect subtle semantic differences that reveal underlying bias.
@@ 6 Conclusion
We contend that FiSCo poses no negative ethical implications for the public; rather, it has the potential for a positive societal impact by identifying subtle bias phenomena within the outputs of LLMs. This capability promotes responsible AI practices, benefiting society as a whole. We collaborated with a professional annotation team to collect labels from human annotators. We provided detailed guidance to the annotators, including an overview of the project, intended use of the labeled data, and the annotation procedure. All annotators were employees located in India and Costa Rica, with equal gender representation (1:1 female to male).
@@
The deployment of LLMs in critical sectors underscores the necessity for comprehensive fairness evaluation frameworks. Our study introduces FiSCo, a fine-grained similarity computation method designed to detect and quantify biases in long-text LLM responses. FiSCo facilitates group-level fairness assessments by decomposing responses into semantically distinct claims and analyzing both intra- and inter-group variances.
Our approach can be used to assess high-stakes applications of LLMs, including hiring, education, and public policy. FiSCo offers a transparent and scalable approach to bias detection. For corporations that leverage LLMs for critical tasks, our method can help identify scenarios where their systems may introduce or reinforce bias.
@@
While FiSCo provides valuable insights into detecting subtle biases within LLMs, it has limitations in analyzing relationships among multiple groups. In our study, when faced with scenarios that involve three or more groups, we conducted pairwise comparisons. This approach limits our ability to draw more nuanced conclusions, such as “the level of bias towards group A is more pronounced than towards group B for group C”. However, in real-world situations, cases involving multiple groups are common, and deriving comprehensive insights beyond pairwise comparisons remains an open challenge.
Furthermore, our evaluation is limited in scope and focuses on a specific set of LLMs, questions, and fairness dimensions. Although this work focuses primarily on developing a fairness evaluation methodology, we acknowledge the need for future research to broaden the scope of analysis to include a wider variety of models and to explore additional dimensions of bias.
Furthermore, bias evaluation goes beyond detecting differences in LLM responses. Being different is a necessary but not sufficient condition for bias. FiSCo does not fully capture the complexity of bias detection, but it serves as an important first step in identifying response differences that may signal bias. It can be further extend to other domains  (Zhang et al., 2025; Xu et al., 2024b; Zhao et al., 2021) and other forms of bias  (Xu et al., 2025; Choi et al., 2024).
Not all differences imply harm, nor do they specify who the bias is directed for or against. The extent or presence of harm in a response often depends on the social context and established stereotypes. For instance, when an LLM suggests a female software engineer to pursue a master’s degree in computer science and a male engineer to pursue a degree in management, the responses may appear neutral on the surface. However, this difference may reflect an underlying bias only when viewed through the lens of societal expectations and stereotypes. Detecting such bias requires social understanding that goes beyond textual analysis.
@@
Appendix A Synthetic Data Generation Details
The prompts for generating Advice and insight-generation questions are shown in Table 6 and Table 7, respectively. However, these initial questions are not directly suitable for our evaluation, as they lack the necessary context. To make them more appropriate, we augment each question with attributes describing the individual. This contextualization ensures the questions reflect realistic user scenarios and allows for meaningful analysis of model responses. We show examples for the two types of questions and responses in Table 8 and Table 9.

### v3

@@ Abstract
Large Language Models (LLMs) often generate responses with inherent biases, undermining their reliability in real-world applications. Existing evaluation methods often overlook biases in long-form responses and the intrinsic variability of LLM outputs. To address these challenges, we propose FiSCo (Fine-grained Semantic Comparison), a novel statistical framework to evaluate group-level fairness in LLMs by detecting subtle semantic differences in long-form responses across demographic groups. Unlike prior work focusing on sentiment or token-level comparisons, FiSCo goes beyond surface-level analysis by operating at the claim level, leveraging entailment checks to assess the consistency of meaning across responses. We decompose model outputs into semantically distinct claims and apply statistical hypothesis testing to compare inter- and intra-group similarities, enabling robust detection of subtle biases. We formalize a new group counterfactual fairness definition and validate FiSCo on both synthetic and human-annotated datasets spanning gender, race, and age. Experiments show that FiSCo more reliably identifies nuanced biases while reducing the impact of stochastic LLM variability, outperforming various evaluation metrics.
@@ 1 Introduction
Several methods have been developed to detect and measure potential biases in LLMs. Demographic representation evaluation methods (Brown et al., 2020; Liang et al., 2022; Mattern et al., 2022; Smith et al., 2022b; Abid et al., 2021; Barikeri et al., 2021) confine the definition of bias to a narrow and commonly used vocabulary. Consequently, these methods fail to capture group-level differences across demographic attributes. Similarly, counterfactual fairness methods (Smith et al., 2022b; Barikeri et al., 2021; Nadeem et al., 2020) typically focus on certain elements in the model’s response rather than the overall response. As a result, they struggle to detect semantic differences in long-form outputs. While some recent efforts (Dwivedi-Yu et al., 2024) recognize the difficulty of capturing subtle biases and attempt to measure LLM fairness from a broader perspective, existing bias benchmarks are limited in length (fewer than 400 characters; see Appendix J for analysis), while GPT-4 produces responses averaging more than 600 characters (Chen et al., 2023).
To address these challenges, we propose FiSCo (Fine-grained Semantic Comparison), a novel statistical framework for evaluating group-level fairness in LLMs by detecting subtle semantic differences in long-form responses across demographic groups (see Figure 1). In contrast to prior methods that rely on predefined bias categories or simplistic text substitutions, our approach captures deeper, group-level disparities in model behavior. Specifically, we extend traditional group fairness concepts to settings where responses are long and variable in structure. FiSCo goes beyond token-level or surface-level comparisons by decomposing each response into semantically distinct claims, and applies state-of-the-art reference checks at the claim level (Hu et al., 2024) to evaluate alignment across responses. We then statistically assess both inter-group and intra-group variability to uncover meaningful differences between demographic groups (see Figure 2). This design enables the detection of subtle biases that would otherwise be obscured by response variability or surface-level text similarity. Furthermore, we leverage statistical hypothesis testing to robustly assess whether observed differences are significant, avoiding the pitfalls of manual prompt manipulation or handcrafted metrics that may overlook real-world patterns. By focusing on semantic content rather than superficial text features, our method enables interpretable and scalable fairness evaluation. To validate the effectiveness of our framework, we construct a comprehensive dataset spanning multiple fairness axes, including gender, race, and age. Experiments demonstrate that FiSCo consistently outperforms existing metrics in detecting nuanced biases with greater stability and alignment to human judgment. The main contributions of this paper are summarized as follows:
@@
  • A new benchmark dataset designed for long-form fairness evaluation, spanning major demographic dimensions (gender, race, and age), enriched with high-quality human annotations and realistic prompt templates, supporting comprehensive, reproducible, and multi-axis fairness assessments across multiple state-of-the-art LLMs.
@@ 2.1 Fairness Evaluation in LLMs
Figure 2: Overview of FiSCo pipeline for evaluating group-level fairness. First, prompts are adapted for each demographic group (e.g., male vs. female), and responses are generated via LLMs. Each response is decomposed into semantic claims, and entailment relationships are computed across response pairs. A fine-grained similarity score is calculated between each response pair. Finally, Welch’s t-test compares inter-group and intra-group similarity distributions to assess the statistical significance of bias.
@@ 2.2 Comparison with Existing Datasets
BBQ (Parrish et al., 2022) Question answering <<20 tokens Race, Gender, Age, etc. Ambiguous QA format
HolisticBias (Smith et al., 2022a) Prompt continuation <<20 tokens 13 social groups Template-based completion
FiSCo (Ours) Long-form generation ~266 chars Gender, Race, Age Open-ended, human-validated
@@ 3.4 Fine-grained Response Similarity
Claim Extraction and Entailment Checking Our approach begins with extracting individual claims from each response. For any two responses $r_{1}$ and $r_{2}$, we first identify the set of claims $C_{1}=\{c_{1}^{(1)},\ldots,c_{1}^{(m)}\}$ from $r_{1}$ and $C_{2}=\{c_{2}^{(1)},\ldots,c_{2}^{(n)}\}$ from $r_{2}$. We then perform a bidirectional semantic entailment check. For each claim $c_{1}^{(i)}$ in $C_{1}$, we determine whether it can be entailed by $r_{2}$, and vice versa for each claim in $C_{2}$ with respect to $r_{1}$. This process results in labeling each claim as one of three categories.
@@
Having multiple claims per response helps our method capture subtle differences between responses. By comparing claims, FiSCo is able to detect fine-grained semantic differences between responses. Similarity Scoring To quantify the similarity between responses, we assign scores in the range of $[0,1]$ to each label type: $\alpha$ for Entailment, $\beta$ for Neutral, and $\gamma$ for Contradiction, where typically $\alpha\geq\beta\geq\gamma$, to reflect that entailment indicates greater similarity than a neutral relationship, which in turn indicates higher similarity than contradiction.
@@
Initially, we set both $\beta$ and $\gamma$ to zero, simplifying our similarity measure to focus exclusively on entailment. This reflects our goal of identifying shared information between responses without penalizing for contradictions or neutrality in the initial analysis phase. For an ablation study varying the value of $\beta$, see Appendix F.1.
@@ 3.5 The Proposed FiSCo Score
With our similarity score function $\mathcal{S}(r_{1},r_{2})$, we can quantitatively assess the fairness of the LLM’s responses across different demographic groups. We do this by comparing inter-group and intra-group similarities. Specifically, for each pair of groups $(G^{\prime},G^{\prime\prime})$, we compute (i) Inter-group similarities: $\mathcal{S}_{\text{inter}}=[\mathcal{S}(r_{i}^{{}^{\prime}},r_{j}^{{}^{\prime\prime}})|i,j\in\{1,\ldots,k\}]$, resulting in $k^{2}$ values, and (ii) Intra-group similarities: $\mathcal{S}_{\text{intra}}=[\mathcal{S}(r_{i}^{{}^{\prime}},r_{j}^{{}^{\prime}})|i,j\in{1,\ldots,k},i<j]+[\mathcal{S}(r_{i}^{{}^{\prime\prime}},r_{j}^{{}^{\prime\prime}})|i,j\in{1,\ldots,k},i<j]$, resulting in k(k1)k(k-1) values.
@@
Using this t-statistic and the degrees of freedom, we calculate the p-value to determine whether the observed differences are statistically significant, with a significance level of $0.05$. Thus, it outputs a binary score. Statistical significance allows FiSCo to avoid inaccurate assessments caused by stochastic variability in LLM responses. FiSCo offers a quantifiable signal of differential treatment in generated outputs across social groups, under controlled conditions. Appendix A provides a High-Level Pseudocode of the FiSCo Algorithm
@@ 4.1 Data Creation
To reflect real-world applicability, each generated template was reviewed by five human annotators via Amazon Mechanical Turk. Annotators were asked to assess whether (1) the question is likely to be asked by a human, and (2) it is suitable for ChatGPT to answer. We retained only those templates for which both questions received a “Yes” with a confidence score above 0.8. Details of our annotation interface, labeling guidelines, and example prompts are included in Appendix C.1.
@@ 4.2 Response Collection
We collect responses from several LLMs, [^1] We use all the LLMs by invoking the Amazon Bedrock APIs (https://aws.amazon.com/bedrock/). More details are given in Appendix K. including Llama 3 70B Instruct (Meta, 2024) and Mixtral 8x7B Instruct (Jiang et al., 2024). The responses were filtered to ensure a minimum length of 30 words. To maintain data integrity, we standardized the prompt structure and balanced demographic categories based on the number of prompts per group. Incomplete or off-topic responses were excluded to ensure that the final dataset contains only high-quality, analyzable outputs. Appendix B includes the prompts and examples used for both question and response generation.
@@ 4.3 Response Evaluation
Details of our SOP, annotation interface, labeling workflows, annotator feedback, and illustrative examples are provided in Appendix C.2. Human-validated templates and fine-grained response comparisons contribute to the realism and reliability of the proposed dataset.
@@ 5 Experiments
For similarity computation, we employ the open-source tool RefChecker (Hu et al., 2024), using Llama 3.1 70B Instruct (Meta, 2024) as both the claim extractor and the checker model (see Appendix H for detailed rationale and Appendix I for ablation study). All experiments were performed on a CPU-based infrastructure, and LLM APIs were accessed through AWS Bedrock. Additional ablation studies are provided in Appendix F.1. In both experiments, we compare the performance of different similarity metrics or group level methods. Here, we use a paired t-test (measuring scores on the same sample across different methods) to assess whether the mean difference in agreement rates between our method and the second-best method (SentenceT5/CBleu) is statistically significant.
@@
5.1 Similarity Metric Evaluation
Experiment Setup. To evaluate the reliability of our similarity measurement Eq. (1), we conducted experiments on two datasets: synthetic and human-annotated. For the synthetic dataset (600 pairs), we generated responses by combining pre-selected claims with pre-defined conditions that specify how the claims should relate. This design enables precise control over ground-truth by explicitly defining claim-level relationships between responses. Details of this process are provided in Appendix B. For the human-annotated dataset (383 pairs), we used the labeled dataset described in Section 4.3. Baseline methods include widely used metrics such as BERTScore and Sentence-BERT, as well as top-performing sentence embedding methods from recent benchmarks (Muennighoff et al., 2022), such as SentenceT5 embeddings.
Baseline Comparison Methods. We compare methods for measuring text similarity, spanning from traditional approaches to state-of-the-art deep learning-based techniques. Bag-of-Words (BoW) (Salton, 1983): Represents text as a collection of word frequencies, without accounting for word order. Term Frequency-Inverse Document Frequency (TF-IDF) (Salton & Buckley, 1988): Assigns higher weights to words that are more distinctive across documents. Word Mover’s Distance (WMD) (Kusner et al., 2015): Measures semantic distance by computing the minimal cumulative distance between word embeddings of two texts. SimCSE (Gao et al., 2021): A contrastive learning-based method that generates sentence embeddings for similarity comparison. BERTScore (Zhang et al., 2019): Uses contextual embeddings from BERT to compute semantic overlap between texts. Sentence-BERT (Reimers, 2019): Extends BERT to produce sentence-level embeddings, enabling efficient similarity computation via cosine distance. SentenceT5 (Ni et al., 2021): A sentence embedding model built on T5, leveraging its text-to-text paradigm to produce high-quality, scalable encodings for various natural language understanding tasks. We adopt largest version of SentenceT5.
@@ 6 Conclusion
We contend that FiSCo does not pose negative ethical implications for the public; rather, it has the potential for a positive societal impact by identifying subtle bias phenomena within the outputs of LLMs. This capability promotes responsible AI practices, benefiting society as a whole. We collaborated with a professional annotation team to collect labels from human annotators. We provide detailed guidance to the annotators, including an overview of the project, the intended use of the labeled data, and the annotation procedure. All annotators were employees located in India and Costa Rica, with equal gender representation (1:1 female to male).
@@
The deployment of LLMs in critical sectors underscores the necessity for comprehensive fairness evaluation frameworks. Our study introduces FiSCo, a fine-grained similarity computation method designed to detect and quantify biases in long-text LLM responses. FiSCo facilitates group-level fairness assessments by decomposing responses into semantically distinct claims and analyzing both intra- and inter-group variances. Our approach can be used to assess high-stakes applications of LLMs, including hiring, education, and public policy. FiSCo offers a transparent and scalable approach to bias detection. For corporations that leverage LLMs for critical tasks, our method can help identify scenarios where their systems may introduce or reinforce bias.
@@
While FiSCo provides valuable insights into detecting subtle biases within LLMs, it has limitations in analyzing relationships among multiple groups. In our study, when faced with scenarios that involve three or more groups, we conducted pairwise comparisons. This approach limits our ability to draw more nuanced conclusions, such as “the level of bias towards group A is more pronounced than towards group B for group C”. However, in real-world situations, cases involving multiple groups are common, and deriving comprehensive insights beyond pairwise comparisons remains an open challenge. Furthermore, our evaluation is limited in scope and focuses on a specific set of LLMs, questions, and fairness dimensions. Although this work focuses primarily on developing a fairness evaluation methodology, we acknowledge the need for future research to broaden the scope of analysis to include a wider variety of models and to explore additional dimensions of bias.
Furthermore, bias evaluation goes beyond detecting differences in LLM responses. Being different is a necessary but not sufficient condition for bias. FiSCo does not fully capture the complexity of bias detection, but it serves as an important first step in identifying response differences that may signal bias. It can be further extend to other domains  (Zhang et al., 2025; Xu et al., 2024b; Zhao et al., 2021) and other forms of bias  (Xu et al., 2025; Choi et al., 2024). Not all differences imply harm, nor do they specify who the bias is directed for or against. The extent or presence of harm in a response often depends on the social context and established stereotypes. For instance, when an LLM suggests a female software engineer to pursue a master’s degree in computer science and a male engineer to pursue a degree in management, the responses may appear neutral on the surface. However, this difference may reflect an underlying bias only when viewed through the lens of societal expectations and stereotypes. Detecting such bias requires social understanding that goes beyond textual analysis.
@@
Appendix A FiSCo Algorithm: High-Level Pseudocode
In this section, we provide a high-level pseudocode description of the FiSCo framework, which quantifies fairness in large language models (LLMs) through semantic and statistical comparisons across demographic groups. FiSCo begins by generating two sets of prompts ($X_{1}$ and $X_{2}$), each tailored to a different demographic group (e.g., gender, race, age), while ensuring semantic equivalence. These prompts are passed through an LLM to obtain two groups of long-form responses ($R_{1}$ and $R_{2}$). Each response is then decomposed into a set of semantically distinct claims. To compare two responses, FiSCo performs bidirectional entailment checks between claims and computes a similarity score based on the proportion of content that is entailed. The algorithm aggregates similarity scores in two ways: intra-group (within-group consistency) and inter-group (across-group consistency). The final fairness score is calculated using Welch’s t-test, which compares the distributions of inter- and intra-group similarities. A statistically significant gap between these distributions suggests a potential bias in the model’s response.
Algorithm 1 FiSCo: Fairness Evaluation via Fine-Grained Semantic Comparison
1:
2:Topic prompt $X$
3:Two demographic groups $G^{\prime}$ and $G^{\prime\prime}$
4:LLM model $M$
5:
6:Fairness score (FiSCo t-statistic and $p$-value)
7:Generate $k$ semantically equivalent prompts for each group: $X_{1}$ for $G^{\prime}$, $X_{2}$ for $G^{\prime\prime}$
8:Generate LLM responses: $R_{1}=M(X_{1})$, $R_{2}=M(X_{2})$
9:for all response pairs $(r_{1},r_{2})$ from $R_{1}$ and $R_{2}$ do
10:  Extract key semantic claims from both responses
11:  Use entailment classification to compare claims between responses
12:  Compute a similarity score $S(r_{1},r_{2})$ based on the proportion of entailed claims
13:end for
14:Compute:
  • Intra-group similarities (within $R_{1}$ and within $R_{2}$)
  • Inter-group similarities (between $R_{1}$ and $R_{2}$)
15:Apply Welch’s t-test to compare inter- and intra-group similarity distributions
16:Return the FiSCo score (t-statistic) and statistical significance ($p$-value)
Appendix B Synthetic Data Generation Details
The prompts for generating Advice and insight-generation questions are shown in Table 5 and Table 6, respectively. However, these initial questions are not directly suitable for our evaluation, as they lack the necessary context. To make them more appropriate, we augment each question with attributes describing the individual. This contextualization ensures that the questions reflect realistic user scenarios and allows for a meaningful analysis of the model responses. We show examples for the two types of questions and responses in Table 7 and Table 8.